# Ikeda type lift on ${\rm SO}(3,n+1)$

### v1

@@ 1. Introduction
Let $\mathbb{A}=\mathbb{A}_{\mathbb{Q}}$ be the ring of adeles of $\mathbb{Q}$. For an integer $n\geq 3$, let $A$ be a positive definite integral matrix of size n2n-2. Assume $A$ is even if $n$ is even while $A={\rm diag}(1,A^{\prime})$ if $n$ is odd where $A^{\prime}$ is a positive definite integral matrix of size n3n-3. Let $J_{2,n}={\rm antidiag}(1,1,-A,1,1)$ and $J_{3,n+1}={\rm antidiag}(1,1,1,-A,1,1,1)$ which are of size $n+3$ and $n+4$ respectively (see Section 2). Let $G={\rm SO}(3,n+1)$ be the special orthogonal group over $\mathbb{Q}$ associated to the symmetric pairing defined by $J_{3,n+1}$. Note that if $n$ is even, $G$ does not have discrete series. Assume $G$ splits everywhere at finite places. By [24, Section 2.1], it is equivalent to $\det(A)=1$ and $n\equiv 2$ mod 8 when $n$ is even, $\det(A^{\prime})=1$ and $n\equiv 3$ mod 8 when $n$ is odd. For instance, this condition is satisfied when $n=8a+2$, $A={\rm diag}(\overbrace{A_{8},\ldots,A_{8}}^{a})$ or $n=8a+3$, $A={\rm diag}(1,\overbrace{A_{8},\ldots,A_{8}}^{a})$ where $A_{8}$ is the $E_{8}$ Cartan matrix given as an element of even integral matrix in $M_{8}(\mathbb{Z})$. This agrees with the result of [8] since $\frac{n+4}{2}\equiv 3$ (mod 4) if $n$ is even, and $\frac{n+3}{2}\equiv 3$ (mod 4).
@@
In [21, Section 4], Pollack computed the Fourier expansion of the Eisenstein series $E_{l}$ on $G(\mathbb{A})$ of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$) along the unipotent radical $N$. It is given by
$E_{l}(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{f}g_{\infty}\in G(\mathbb{A})$
where $\mathcal{W}_{2\pi\eta}$ is Pollack’s spherical function associated to the additive character $e^{2\pi\sqrt{-1}(\eta,\ast)_{V^{\prime}}}$ on $V^{\prime}(\mathbb{R})$ (see [21, Section 3] but we use a slightly different normalization as explained later). We remark that $a_{E_{l}}(\eta)=0$ unless $\eta\in V^{\prime}(\mathbb{Z})$.
@@
Keep the notation being as above. Let $l>n+1$ be an even integer. Let $f$ be a Hecke eigen cusp form of weight $k:=l-\frac{n-2}{2}$ with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $f(z)=\displaystyle\sum_{n=1}^{\infty}a_{f}(n)n^{\frac{k-1}{2}}q^{n}$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $A_{f}(\eta)(g_{\mathbb{f}})=q(\eta)^{\frac{k-1}{2}}\displaystyle\prod_{p}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(\alpha_{p})$ for $g_{\mathbb{f}}=(g_{p})_{p}\in G(\mathbb{A}_{\mathbb{f}})$. Then
@@
is a Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F_{f}}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
@@
When $n$ is odd, we need a slight modification. Let $l>n+1$ be an even integer. Let $f$ be a Hecke eigen cusp form of weight k:=2ln+1k:=2l-n+1 with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $f(z)=\displaystyle\sum_{n=1}^{\infty}a_{f}(n)n^{\frac{k-1}{2}}q^{n}$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $h(z)=\displaystyle\sum_{n>0\atop\text{$\epsilon n\equiv 0,1$ (mod 4)}}c(n)q^{n}\in S_{l-\frac{n}{2}+1}(\Gamma_{0}(4))$ be the one corresponding to $f$ by the Shimura correspondence, where $\epsilon=(-1)^{l-\frac{n-1}{2}}$. Write $q(\eta)=\mathfrak{d}_{\eta}\mathfrak{f}_{\eta}^{2}$ such that $\mathfrak{d}_{\eta}$ is the absolute discriminant of $\mathbb{Q}(\sqrt{\pm q(\eta)})/\mathbb{Q}$, where the sign is such that $\pm q(\eta)\equiv 0,1$ (mod 4). Let $\chi_{\eta}$ be the primitive Dirichlet character corresponding to $\mathbb{Q}(\sqrt{\pm q(\eta)})/\mathbb{Q}$. Then for each rational prime $p$, there exists a Laurent polynomial $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p})=\widetilde{Q}_{\eta,p}(X_{p})\in\mathbb{Z}[X_{p},X^{-1}_{p}]$ depending on $\eta$ and $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}$ satisfying $\widetilde{Q}_{\eta,p}(X_{p})=\widetilde{Q}_{\eta,p}(X^{-1}_{p})$ such that
@@
is a Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
@@
We organize this paper as follows. In Section 2, we give explicit forms of $P$ and another maximal parabolic $Q$ which gives rise to the Jacobi group $J$ in our setting. In Section 3, we recall Pollack’s definition of certain types of automorphic forms on $G(\mathbb{A})$ [21]. One of the important ingredients is an explicit description of the Siegel series and using this, we deduce the functional equation for $\widetilde{Q}_{\eta,p}$ which plays an important role in our construction. The computation of the Siegel series is complicated and it is contained in Section 11. The reader may assume the results therein to understand the proof of the main theorems. One corollary of our explicit computation is that we can show that Fourier coefficients of Eisenstein series have bounded denominator. (See Theorem 4.9 and 4.10.) In [21], Pollack can only show algebraicity of the Fourier coeffcients. In Section 5, we briefly recall the Weil representation for $J$. The Fourier-Jacobi expansion along $J$ of Eisenstein series associated to $P$ fixed by G(^)G($\widehat{\mathbb{Z}}$) is also explicitly computed in Section 6 and 7. Then, applying the theory of a compatible family of Eisenstein series due to Ikeda [10], [11], we give a uniform proof for the main theorems in Section 8. We also compute the $L$-functions of our lifts in Section 9, and interpret our lifts in terms of CAP representations in Section 10. Finally, Appendix is given for computing some integrals which show up in Section 7.
@@ 2. Preliminaries
so that $(e,f)_{V}=eJ_{3,n+1}{}^{t}f=1$. Let $q$ be the quadratic form on $V^{\prime}$ defined by $J_{2,n}$. Denote by
@@
for $x,y\in V^{\prime}$ the associated bilinear form. The quadratic form on $V$ is $\tilde{q}(\alpha e+v^{\prime}+\beta f)=\alpha\beta+q(v^{\prime})$.
@@ 2.1. Siegel parabolic subgroup $P=MN$
Then, ${\rm diag}(t,m,t^{-1})n\in P$ acts on $x\in V^{\prime\prime}$ by $t^{2}(m\cdot x)$ where $m\cdot x$ stands for the usual matrix multiplication.
@@ 3. Automorphic forms on $G={\rm SO}(3,n+1)$
By definition, vectors $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$, are anisotropic vectors and called rank 2 elements in [21].
@@ 4. Eisenstein series
In this section, we follow [21, Section 4]. For any $l>n+1$ and a Schwartz-function $\Phi_{\mathbb{f}}$ on $V(\mathbb{A}_{\bf f})$, we define a section $f$ of ${\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}\,|\nu|^{s-\frac{n+2}{2}}$ (normalized induction) by
@@
where $\mathcal{W}_{2\pi\eta}(g_{\infty})$ is the Pollack’s spherical function. Let $\Phi_{\mathbb{f}}=\otimes_{p}\Phi_{p}\otimes\Phi_{\infty}$, and $\Phi_{p}$ is unramified for all $p$. We compute explicitly $a_{E_{l}}(\eta)$. If $\eta$ is of rank 2, i.e., $q(\eta)\neq 0$,
(4.2) $a_{E_{l}}(\eta)(g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty})=J(s,\eta,f_{l})=\int_{V^{\prime}(\mathbb{A})}\psi((\eta,x))f_{l}(wn(x),g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}},s)\,dx.$
@@
$J(s,\eta,f_{l})=\prod_{p}J(s,\eta,\Phi_{p}),\ J(s,\eta,\Phi_{p}):=\int_{V^{\prime}(\mathbb{Q}_{p})}\psi_{p}((\eta,x))f_{l}(wn(x),g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}},s)\,dx.$
Notice that $J(s,\eta,\Phi_{p})$ is also defined for $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$ and it is easy to see that $J(s,\eta,\Phi_{p})=0$ if $\eta\not\in V^{\prime}(\mathbb{Z}_{p})$. The following lemma is an analogue of [16, p.591, Lemma 3.4] and it is proved by the same argument therein:
@@ 4.1. $p$-adic computation: Siegel series
$B_{r,\eta}=\int_{p^{-r}V^{\prime}(\mathbb{Z}_{p})}\psi((\eta,x))\text{char}(p^{r}q(x)\in\mathbb{Z})\,dx=\int_{V^{\prime}(\mathbb{Z}_{p})\atop q(x)\equiv 0\,\text{mod $p^{r}$}}\psi\Big(\frac{(\eta,x)}{p^{r}}\Big)\,dx.$
Set $x=u+y$, where $u\in V^{\prime}(\mathbb{Z}_{p}/p^{r}\mathbb{Z}_{p})$, and $y\in V^{\prime}(p^{r}\mathbb{Z}_{p})$. If $y\equiv 0$ (mod $p^{r}$), $q(y)\equiv 0$ (mod $p^{r}$). Hence
@@
Suppose $n$ is even. Let $p^{k}\|\eta$, and $p^{2k+k^{\prime}}\|q(\eta)$. Then $J(s,\eta,\Phi_{p})=(1-p^{m-1-s})Q_{\eta,p}(p^{m-s})$, where $Q_{\eta,p}(X)\in\mathbb{Z}[X]$ is a polynomial of degree $2k+k^{\prime}$, and it satisfies the functional equation
@@
Remark 4.4.
Here $\chi_{\eta}$ is the primitive Dirichlet character corresponding to $\mathbb{Q}(\sqrt{\pm q(\eta)})/\mathbb{Q}$. Write $q(\eta)=\mathfrak{d}_{\eta}\mathfrak{f}_{\eta}^{2}$ such that $\mathfrak{d}_{\eta}$ is the discriminant of $\mathbb{Q}(\sqrt{\pm q(\eta)})/\mathbb{Q}$. If $2k+k^{\prime}$ is odd, then $p|\mathfrak{d}_{\eta}$ and hence $\chi_{\eta}(p)=0$.
@@
Let $n$ be odd and $p=2$. Let $2^{k}\|\eta$ and $2^{2k+k^{\prime}}\|q(\eta)$. Let $l=\lfloor\frac{2k+k^{\prime}}{2}\rfloor$. Let $\chi_{\eta}$ be the quadratic character associated to $\mathbb{Q}(\sqrt{\pm q(\eta)})/\mathbb{Q}$. If $q(\eta)$ is a perfect square, let $\chi_{\eta}(2)=1$. If 2 ramifies in $\mathbb{Q}(\sqrt{\pm q(\eta)})$, $\chi_{\eta}(2)=0$. Then
@@
Here if $k^{\prime}$ is odd, then $\chi_{\eta}(2)=0$. If $k^{\prime}$ is even, and $q(\eta)2^{-2k-k^{\prime}}\equiv 3$ (mod 4), then $\chi_{\eta}(2)=0$. If $k^{\prime}$ is even, and $q(\eta)2^{-2k-k^{\prime}}\equiv 1$ (mod 4), then $\chi_{\eta}(2)=(\tfrac{q(\eta)2^{-2k-k^{\prime}}}{2})$, where for $n\equiv 1$ (mod 4), $(\tfrac{n}{2})=\begin{cases}1,&\text{if $n\equiv 1$ (mod 8)}\\ -1,&\text{if $n\equiv 5$ (mod 8)}\end{cases}$.

### v2

@@ 1. Introduction
Let $\mathbb{A}=\mathbb{A}_{\mathbb{Q}}$ be the ring of adeles of $\mathbb{Q}$. For an integer $n\geq 3$, let $A$ be a positive definite integral matrix of size n2n-2. Assume $A$ is even (namely, any diagonal entries are even integers) if $n$ is even while $A={\rm diag}(1,A^{\prime})$ if $n$ is odd where $A^{\prime}$ is a positive definite integral matrix of size n3n-3. Let $J_{2,n}={\rm antidiag}(1,1,-A,1,1)$ and $J_{3,n+1}={\rm antidiag}(1,1,1,-A,1,1,1)$ which are of size $n+3$ and $n+4$ respectively (see Section 2). Let $G={\rm SO}(3,n+1)$ be the special orthogonal group over $\mathbb{Q}$ associated to the symmetric pairing defined by $J_{3,n+1}$. Note that if $n$ is even, $G$ does not have discrete series. Assume $G$ splits everywhere at finite places. By [26, Section 2.1], it is equivalent to $\det(A)=1$ and $n\equiv 2$ mod 8 when $n$ is even, $\det(A^{\prime})=1$ and $n\equiv 3$ mod 8 when $n$ is odd. For instance, this condition is satisfied when $n=8a+2$, $A={\rm diag}(\overbrace{A_{8},\ldots,A_{8}}^{a})$ or $n=8a+3$, $A={\rm diag}(1,\overbrace{A_{8},\ldots,A_{8}}^{a})$ where $A_{8}$ is the $E_{8}$ Cartan matrix given as an element of even integral matrix in $M_{8}(\mathbb{Z})$. This agrees with the result of [8] since $\frac{n+4}{2}\equiv 3$ (mod 4) if $n$ is even, and $\frac{n+3}{2}\equiv 3$ (mod 4).
@@
In [23, Section 4], Pollack defined the Eisenstein series $E_{l}$ on $G(\mathbb{A})$ of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$), and computed unramified and archimedean part of Fourier coefficients. He can show the algebraicity of the Fourier coefficients. However, he did not compute the Siegel series explicitly which come from the bad finite places. The computation of the Siegel series is complicated and it is contained in Section 11. One corollary of our explicit computation is that we can show that Fourier coefficients of Eisenstein series have bounded denominator. (See Theorem 4.8 and 4.9.) After a long computation, we can write down the Fourier expansion of the Eisenstein series as follows:
$E_{l}(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=\gamma k_{f}g_{\infty}\in G(\mathbb{A})=G(\mathbb{Q})(G(\widehat{\mathbb{Z}})\times G(\mathbb{R}))$
where $\mathcal{W}_{2\pi\eta}$ is Pollack’s spherical function associated to the additive character $e^{2\pi\sqrt{-1}(\eta,\ast)_{V^{\prime}}}$ on $V^{\prime}(\mathbb{R})$ (see [23, Section 3] but we use a slightly different normalization as explained later). We remark that $a_{E_{l}}(\eta)=0$ unless $\eta\in V^{\prime}(\mathbb{Z})$.
@@
Keep the notation being as above. Let $l>n+1$ be an even integer. Let $f$ be a normalized Hecke eigen cusp form of weight $k:=l-\frac{n-2}{2}$ with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $f(z)=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $A_{f}(\eta)(g_{\mathbb{f}})=q(\eta)^{\frac{k-1}{2}}\displaystyle\prod_{p}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(\alpha_{p})$ for $g_{\mathbb{f}}=(g_{p})_{p}\in G(\mathbb{A}_{\mathbb{f}})$. Then
@@
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F_{f}}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
@@
When $n$ is odd, we need a slight modification. Let $l>n+1$ be an even integer. Let $f$ be a normalized Hecke eigen cusp form of weight k:=2ln+1k:=2l-n+1 with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $f(z)=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $h(z)=\displaystyle\sum_{m>0\atop\text{$(-1)^{\frac{k}{2}}m\equiv 0,1$ (mod 4)}}c(m)q^{m}\in S_{l-\frac{n}{2}+1}(\Gamma_{0}(4))$ be the one corresponding to $f$ by the Shimura correspondence. Write $q(\eta)=\mathfrak{d}_{\eta}\mathfrak{f}_{\eta}^{2}$ such that $\mathfrak{d}_{\eta}$ is the absolute discriminant of $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$, where $\epsilon=\begin{cases}1,&\text{if $q(\eta)_{1}\equiv 1$ (mod 4)}\\ -1,&\text{if $q(\eta)_{1}\equiv 3$ (mod 4)}\end{cases}$, and $q(\eta)=2^{v_{2}(\eta)}q(\eta)_{1}$, $q(\eta)_{1}$ odd. Let $\chi_{\eta}$ be the primitive Dirichlet character corresponding to $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$. Then for each rational prime $p$, there exists a Laurent polynomial $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p})=\widetilde{Q}_{\eta,p}(X_{p})\in\mathbb{Z}[X_{p},X^{-1}_{p}]$ depending on $\eta$ and $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}$ satisfying $\widetilde{Q}_{\eta,p}(X_{p})=\widetilde{Q}_{\eta,p}(X^{-1}_{p})$ such that
@@
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
@@
We organize this paper as follows. In Section 2, we give explicit forms of $P$ and another maximal parabolic $Q$ which gives rise to the Jacobi group $J$ in our setting. In Section 3, we recall Pollack’s definition of certain types of automorphic forms on $G(\mathbb{A})$ [23]. In Section 4, we recall the definition of Eisenstein series, and by assuming the explicit calculation of the Siegel series in Section 11, we write down the formula for the Fourier coefficients of the Eisenstein series. In Section 5, we briefly recall the Weil representation for $J$. The Fourier-Jacobi expansion along $J$ of Eisenstein series associated to $P$ fixed by G(^)G($\widehat{\mathbb{Z}}$) is also explicitly computed in Section 6 and 7. Then, applying the theory of a compatible family of Eisenstein series due to Ikeda [10], [11], we give a uniform proof for the main theorems in Section 8. We also compute the $L$-functions of our lifts in Section 9, and interpret our lifts in terms of CAP representations in Section 10. Finally, Appendix is given for computing some integrals which show up in Section 7.
@@ 2. Preliminaries
so that $(e,f)_{V}=\frac{1}{2}eJ_{3,n+1}{}^{t}f=\frac{1}{2}$. Let $q$ be the quadratic form on $V^{\prime}$ defined by $q(x)=\frac{1}{2}xJ_{2,n}{}^{t}x$ for $x\in V^{\prime}$. Denote by
@@
for $x,y\in V^{\prime}$ the associated bilinear form. The quadratic form $\tilde{q}$ on $V$ for $J_{3,n+1}$ satisfies $\tilde{q}(\alpha e+v^{\prime}+\beta f)=\alpha\beta+q(v^{\prime})$.
@@ 2.1. Siegel parabolic subgroup $P=MN$
Then, ${\rm diag}(t,m,t^{-1})n\in P$ acts on $x\in V^{\prime}$ by $t^{2}(m\cdot x)$ where $m\cdot x$ stands for the usual matrix multiplication.
@@ 3. Automorphic forms on $G={\rm SO}(3,n+1)$
By definition, vectors $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$ are anisotropic vectors and they are called rank 2 elements in [23].
@@ 4. Eisenstein series
In this section, we follow [23, Section 4]. For any $l>n+1$ and a Schwartz-function $\Phi_{\mathbb{f}}$ on $V(\mathbb{A}_{\bf f})$, we define a section $f$ of ${\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}\,|\nu|^{s-\frac{n+2}{2}}$ (normalized induction) by
@@
[…]
(4.2) $E_{l}(g,\Phi_{\mathbb{f}})=E_{0}(g)+\sum_{0\neq\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)(\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty})$
for $g=\gamma k_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})=G(\mathbb{Q})(G(\widehat{\mathbb{Z}})\times G(\mathbb{R}))$, since $k_{\mathbb{f}}\cdot\Phi_{\mathbb{f}}=\Phi_{\mathbb{f}}$.
We compute explicitly $a_{E_{l}}(\eta):=a_{E_{l}}(\eta)(\Phi_{\mathbb{f}})$. If $\eta$ is of rank 2, i.e., $q(\eta)\neq 0$,
(4.3) $a_{E_{l}}(\eta)(\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty})=J(s,\eta,f_{l})=\int_{V^{\prime}(\mathbb{A})}\psi((\eta,x))f_{l}(wn(x),\Phi_{\mathbb{f}},s)\,dx.$
@@
$J(s,\eta,f_{l})=\prod_{p}J(s,\eta,\Phi_{p}),\ J(s,\eta,\Phi_{p}):=\int_{V^{\prime}(\mathbb{Q}_{p})}\psi_{p}((\eta,x))f_{l}(wn(x),\Phi_{\mathbb{f}},s)\,dx$
where $w\in G={\rm SO}(V)$ is defined by $w\cdot e=f$, $w\cdot f=e$, and $w|_{V^{\prime}}={\rm id}_{V^{\prime}}$. Notice that $J(s,\eta,\Phi_{p})$ is also defined for $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$. Further, by [23, p.633, line -6], $J(s,\eta,\Phi_{p})=\displaystyle\int_{{\rm GL}_{1}(\mathbb{Q}_{p})}\int_{V^{\prime}(\mathbb{Q}_{p})}\psi_{p}((\eta,x))\Phi_{p}(t,tx,-tq^{\prime}(x)))|t|^{s}_{p}\,dxdt$. Since $t$ runs over $\mathbb{Z}^{\times}_{p}$, $\Phi_{p}(t,tx,-tq^{\prime}(x)))$ is preserved under the transformation $x\mapsto x+x_{0}$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. By the change of variables, $J(s,\eta,\Phi_{p})=\psi((\eta,x_{0}))J(s,\eta,\Phi_{p})$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. Thus, $J(s,\eta,\Phi_{p})=0$ if $\eta\not\in V^{\prime}(\mathbb{Z}_{p})$.
The following lemma is an analogue of [16, p.591, Lemma 3.4] and it is proved by the same argument therein:
@@ 4.1. $p$-adic computation: Siegel series
$B_{r,\eta}=\int_{p^{-r}V^{\prime}(\mathbb{Z}_{p})}\psi((\eta,x))\text{char}(p^{r}q(x)\in\mathbb{Z})\,dx=p^{r(n+2)}\int_{V^{\prime}(\mathbb{Z}_{p})\atop q(x)\equiv 0\,\text{mod $p^{r}$}}\psi\Big(\frac{(\eta,x)}{p^{r}}\Big)\,dx.$
Set $x=u+y$, where $u\in V^{\prime}(\mathbb{Z}_{p}/p^{r}\mathbb{Z}_{p})$, and $y\in V^{\prime}(p^{r}\mathbb{Z}_{p})$. If $y\equiv 0$ (mod $p^{r}$), $q(y)\equiv 0$ (mod $p^{r}$). It follows from this that, upon making the change of variables in $y$, the factor $p^{-r(n+2)}$ arises, cancelling the factor $p^{r(n+2)}$ obtained earlier. Thus
@@
Suppose $n$ is even. Let $p^{k}\|\eta$, and $p^{2k+k^{\prime}}\|q(\eta)$. Then
$J(s,\eta,\Phi_{p})=(1-p^{m-1-s})Q_{\eta,p}(p^{m-s}),$
where $Q_{\eta,p}(X)\in\mathbb{Z}[X]$ is a polynomial of degree $2k+k^{\prime}$, and it satisfies the functional equation
@@
Recall that $(\tfrac{n}{2})=0$ if $2|n$ or $n\equiv 3$ (mod 4). For $n\equiv 1$ (mod 4), $(\tfrac{n}{2})=\begin{cases}1,&\text{if $n\equiv 1$ (mod 8)}\\ -1,&\text{if $n\equiv 5$ (mod 8)}\end{cases}$.
@@
Let $n$ be odd and $p=2$. Let $2^{k}\|\eta$ and $2^{2k+k^{\prime}}\|q(\eta)$. Let $l=\lfloor\frac{2k+k^{\prime}}{2}\rfloor$. Let $q(\eta)_{1}=q(\eta)2^{-2k-k^{\prime}}$. Let $\chi_{\eta}(2)=\begin{cases}(\frac{q(\eta)_{1}}{2}),&\text{if $2k+k^{\prime}$ even}\\ 0,&\text{if $2k+k^{\prime}$ odd}\end{cases}$. Then